OpenStudy (kawii2004):

Which algebraic expression is equivalent to the expression below? 7(3x + 6) + 19 63x 21x + 61 7x + 9 21x + 23

jimthompson5910 (jim_thompson5910):

first distribute 7(3x + 6) = 7*3x + 7*6 = ??? notice how I multiplied the outer 7 by each term inside the parenthesis

OpenStudy (jamesk):

And don't forget after you have distributed 7(3x + 6) into 7*3x + 7*6 as suggested earlier, that you also need to add 19 to the result, yielding 7*3x + 7*6 +19